Prev Next

SW&IT Meeting Minutes 2018-10-08

Attendees: Andrew, David, Igor, Kellee, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

News from management

  • EM #3 interview
  • board meeting

Previous Action Items

  • Vsphere (DE)
  • power plan update draft (DE)
  • Luci telemetry
    • web-io cold heads state for vibration (IS)
    • luci telemetry in git (DONE)
  • vm backup

Transition plan

Handover steps

  1. primary + secondary do self learning, gather information, read wiki
  2. primary organize teamviewer/zoom walk through with Kellee and secondary
  3. primary draft techtalk w. input from secondary
  4. techtalk review by Kellee
  5. primary does a team presentation

Please update this table to indicate progress.


Application(s)Sorted ascending Primary Secondary Week
ALH Igor(1/2)
Matthieu (2)
1
AzCam
AC (2)
Petr (2)
2
DIMM
AC (4)
Petr (2)
2
FACSUM,WEATHERSTATION
AC (2)
Matthieu (2)
4
GCS
Petr (3)
Matthieu (2)
1
INDI Igor Matthieu 3
LBC
Yang (3)
Igor(1/2) 2
MODS
Yang (2)
Igor 2
OAC
AC (4)
Petr (4)
1
OVMS
Petr (3)
Matthieu (2)
3
particle monitors AC Petr 4
Status Server Stephen (2)
Matthieu (2)
2
Telem-VIZ
Yang (1)
Stephen (1) 3
Telemetry
Petr (3)
Matthieu (2)
2
Telemetry cleanup Stephen (1) Petr 3
Weatherstation
AC (2)
Petr (2)
2

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • FOS Wiki transition
  • Magic laptop clone/vm (news from Microgate?)
  • SOUL/FLAO Oct18 in lab then telescope
  • LMS tcl/tk?
  • INDI / AO

This week in github

Team reports

PK

  • AGw4 support, work
  • cleaned old unused variables from OAC
  • reviewed AC AOC HO presentation
  • added set foldmethod=syntax, set foldstart=99 in .vimrc
  • find a possible cause for IT7387 (LBC NS tracking), working on tests
  • modified pcsclient to support non interactive testing
  • dust cleaned office monitors and stands
next:

SH

  • Foswiki final cleanup from problem reports
  • Bridge between networks (Mountain-in-Lab environment).
  • Fedora 28 burn for Xianyu and I.
  • AGW networking in clean room.
  • Docker cluster work (figured out how to assign static IPs to containers, added extra storage, NFS volumes, etc.), docker registry work.
  • Request changes to DHCP from to add UEFI PXE boot as well as BIOS PXE boot.
  • Request to add docker-registry.lbto.org to DNS.
  • AGW3 in cleanroom.
  • Rebuild LAMP server for Al Conrad on Fedora, Maria, etc.
Next:
  • Docker registry finalization.
  • Convert RPM builds in Jenkins to follow new recipe.
  • Experiment with passing hardware resources through to containers.
  • Light Foswiki?
  • AGW3 in cleanroom.

YZ

  • LBC and MODS handover:
    • Starting to prepare LBC tech talk document.
    • Continue to read MODS wiki documents.
  • Al's AO-CRAVAT project:
    • Met with Al for a brief discussion, we agreed to follow Al's previous project proposal concept, but that the details (such as what database to use, programming language/components/libraries are left for me to fill in).
    • After discussion with Matthieu, we decided for now to use a relational DB for data storage and query (going to use sqlite for now), and Google's Polymer.js for web front end.
    • The whole architecture is yet to be determined, but we are leaning to use Polymer.js for web front, Apache for web server to handle communication, and Python/C++ etc. as background process that does the work. There is also a plan to integrate command-line interface in addition to the web front interface.
    • Learning Javascript and Polymer.js
    • This is also an effort to propose and create a standard architecture and workflow for much of the other similar work in LBTO (e.g., telemetry query/visualization, etc.)
  • Other continuing stuffs such as Python tools, etc.

MB

  • Luci H/O doc sent to Mark W.
  • shark-vis discussions
  • jenkins-ci rework
  • argos deploy with AC and github/mpia follow up
  • eng mgr 3 interview
  • ims kicktstart + online, notice on swgdev01 old site
  • polymer 3
  • PBP kickstart meeting
  • review AO cravat new document
  • OVMS+TT meeting, discuss AO cabinet with Michael W.
next:
  • sci-arch meeting prep
  • Luci H/O doc in VCAN
  • startech LBC update
  • review Guido PBP requirements document

AC

Last week
  • Particle monitor telemetry applications
  • AGW UMAC telemetry
  • OAC tech talk reviewed by AC and PK
  • ARGOS deployment
  • Vacation Wednesday and Thursday

This week:
  • ARGOS deployment
    • github issues #1 and #3
  • Tech talks to be scheduled
  • Vacation Tuesday morning

KS

Last week
  • Updated the DIMM User's Guide in the Vault with auto-acq info.
  • Updated particle monitor wiki pages - added CDA, contact info, Vault doc references.
  • Looked into strange power down behavior on LBC that Olga saw using "Warm Up".
  • Reviewed most of my hardcopy binders for notes to pass on.
  • Started cleaning up my home directory in Tucson.
  • Reviewed OAC tech talk and sent notes to AC and PK.
  • Vacation day on Thursday.

This week
  • Continue going through all my application notes, binders, wiki pages, email to pass on relevant info.

XZ

Last week
  • Try to install Engpanel (currently running on the AO laptop) on Window 7 and 10, both not work
  • check the swith BCU firmware version on both sides on the telescope
  • Test the power back plane lab setup with mini-coverter (works)
  • send email to Guido, which send MG questions aboout the Engpanel and BCU firmware
  • AGW connection test
  • Compile the AO INDI on simulator and check in more necessary files to SVN
  • Combine the AO monitor scripts for SX and DX (tested and installed)
  • Power backplane meeting
This week
  • read power backplane requirement
  • support the on-going AO work

DE

Last week:
  • Talked to Olga re: LBT wireless in 507
    • No recent complaints of the AP not working or being insufficient for the teams' work
    • Last problem requiring reboot was in January
  • Backups to GDrive
    • Making progress working around Google upload limits and avoiding saturation of the Synology's network connection
  • Progress on server room power proposal
    • Have meeting this afternoon with Liebert rep
  • Started troubleshooting ailing Perle network switch, aiming for replacement.
  • Test restore of selected VMs
    • As far as I am able to tell, our current restores are fully functional machines using the VMWare Tools quiescence.
This week:
  • Backups monitoring
  • Perle switch
  • Jury duty redux Tuesday +

IS

Last week

  • LUCI
    • Client application for LUCI don't work over VPN. I'm not clear of a solution for this, as still don't understand the problem fully. This is not an issue but it could make life easier, particularly for support.
    • Development of RTD improvements and bug fix.
      • The development, (I think) is done, I need to fully test it. That means generating a fits image with a MOS mask in it and expect to see the new headers, and have the RTD updated with the new image and LMS information, repeat this with another image with a different mask.
  • OCS
    • OT
      • Mixmode operation, MODS/LBC in binocular mode. Michelle E. brought this issue last week but couldn't give it much time, probably will be something we'll look at this week.
  • PIT
    • Added PDF generation on pit.lbto.org for 2019A.1.0.0, and pending on Michelle decision, will install 2019A.1.2.0 probably this week.

This week:

  • INDI/ALH/LUCI/MODS
    • I need to get back on this and really give them to good push.
  • LUCI
    • Wrapping up RTD improvements.
    • Testing LUCI
    • Will take the opportunity to learn a bit more about the sequencer/journalizer/GEIRS connection in LUCI.
    • Write a small program that generates, reads/consumes LUCI journalizer events
  • OT
    • Possibly mixmode testing/research
    • Possibly start with AFC changes if times allow.
Topic revision: r15 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback